OpenStudy (anonymous):

Translate the sentence into an equation. Four times the difference of a number and 9 equals 7. Use the variable x for the unknown number Will medal

Vocaloid (vocaloid):

any ideas?

OpenStudy (anonymous):

no

OpenStudy (kittiwitti1):

Four times: x4 Difference of number c and 9: c-9 Equals 7: =7 Put together:\[4\times(c-9)=7\]I may have gotten the difference part wrong. @Vocaloid is it c-9 or 9-c?

OpenStudy (anonymous):

ty